OpenStudy (anonymous):

MEDALS A fossilized leaf contains 15% of its normal amount of carbon 14. How old is the fossil (to the nearest year)? Use 5600 years as the half-life of carbon 14. Solve the problem. what would be the number??

OpenStudy (anonymous):

first write 15% as \(.15\) because math uses numbers, not percents then to right to the equation \[\large \left(\frac{1}{2}\right)^{\frac{t}{5600}}=.15\] solve for \(t\) in one step (plus a calculator)

OpenStudy (anonymous):

you know how to solve it ? if not, let me know it actually takes two steps, i lied

OpenStudy (anonymous):

CAnt figure it out on my calc?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what are you trying to enter?

OpenStudy (anonymous):

the whole equation

OpenStudy (anonymous):

lol no you cannot enter the whole equation, you have to solve for \(t\)

OpenStudy (anonymous):

lets go slow first of all, is it clear where i got \[\large \left(\frac{1}{2}\right)^{\frac{t}{5600}}=.15\] from ?

OpenStudy (anonymous):

im 23 doing 12th grade math ... help me out lol

OpenStudy (anonymous):

online thing

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

ok lets first of all get the answer, and then see where the equation came from

OpenStudy (anonymous):

wud it b 15299

OpenStudy (anonymous):

i hvae 3 minutes to put my answer in

OpenStudy (anonymous):

the way you solve \[b^x=A\] for \(x\) is by the "change of base" formula \[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] in this example \(x=\frac{t}{5600}, b=\frac{1}{2}=.5, A=.15\) and so \[\large \left(\frac{1}{2}\right)^{\frac{t}{5600}}=.15\] \[\large \iff \frac{t}{5600}=\frac{\ln(.15)}{\ln(.5)}\]making \[t=\frac{5600\ln(.15)}{\ln(.5)}\]

OpenStudy (anonymous):

i get \(15327\)

OpenStudy (anonymous):

not correct not one of my answers...

OpenStudy (anonymous):

pick the one closest then

OpenStudy (anonymous):

well that means the problem wasnt done correctly but ok

OpenStudy (anonymous):

no actually i did it two different ways and arrived at the same answer it means your answers are not correct, or someone did it not carefully enough

OpenStudy (anonymous):

what is the answer closest to \(15327\) ?

OpenStudy (anonymous):

15299 and 15327 are only 28 years apart, which is not much in the scale of 15 thousand years i bet the grad student who wrote this problem rounded somewhere